# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

### v2

@@ 3.1   Estimation of $\beta$-mixing coefficients under general ergodicity condition
We use the shortcut notation $N_{x}=N_{x}^{(1)}$ and $N_{xx^{\prime}}=N_{xx^{\prime}}^{(1)}$ to respectively denote number of visits to state $x$ and transitions from $x$ to $x^{\prime}$ on the original chain. We define the following estimator $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ for $\beta(s)$,
@@
$\widehat{\beta}(s)=\frac{1}{2\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|,$ (5)
and by convention, we set $\widehat{\beta}(s)=0$ for $s\geq n$.
Theorem 3.1.
@@
Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, it holds that
@@
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{\mathcal{B}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}}\left(\sqrt{\frac{3}{2}\mathcal{J}_{p}^{(s)}}+2\sqrt{\left\|\pi\right\|_{1-1/p}^{1-1/p}}\right)\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
where for $x,x^{\prime}\in\mathcal{X}$, we wrote $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$, and
@@
$\mathcal{B}_{p}^{(s)}\doteq\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p},\qquad\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}.$
We observe that the two key quantities $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$ control the estimation rate. The former essentially captures $\beta$-mixing properties of the process, while the latter is tightly connected to entropic properties of the distribution of (skipped) stationary pairs. In the following Corollary 3.1, we readily provide more explicit upper bounds on $\mathcal{B}_{p}^{(s)}$ under additional assumptions on the mixing rate. The analysis of $\mathcal{J}_{p}^{(s)}$ is deferred to Section 4.
Corollary 3.1.
@@
Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$ with $\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}<\infty$, the following statements hold.
@@
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}\left(\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}\right)\mathcal{J}_{p}^{(s)}}.\\ \end{split}$
@@
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{\beta_{0}^{1/p}}{\lfloor(n-1)/s\rfloor}\left(1+\exp(-\beta_{1}s^{b}/p)+\frac{\Gamma\left(\frac{1}{b},s^{b}\beta_{1}/p\right)}{bs(\beta_{1}/p)^{1/b}}\right)\mathcal{J}_{p}^{(s)}},\end{split}$
[…]
@@
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{\beta_{0}^{1/p}+\zeta(b/p)\frac{\beta_{1}^{1/p}}{s^{b/p}}}{\lfloor(n-1)/s\rfloor}\mathcal{J}_{p}^{(s)}},\end{split}$
@@
3.2   Comparison with prior results
In McDonald et al. (2015, Theorem 4), the authors establish the nearly parametric convergence rate for continuously valued Markov processes
where $W_{0}$ is the Lambert function (Corless et al., 1996). They also posit that their findings may be considered conservative when applied to discrete-valued processes. Theorem 3.1 validates their remark by recovering the superior parametric rate $\mathcal{O}(\sqrt{1/n})$ under convergence of $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$. Our method shares similarities with Khaleghi and Lugosi (2023), as both employ Rio’s covariance inequality Rio (1999, Corollary 1.1). However, compared to both Khaleghi and Lugosi (2023) and McDonald et al. (2015, Theorem 4), we distinguish ourselves by focusing on the discrete space setting and deriving explicit rates in terms of fine-grained properties of the transition operator.
3.3   Estimation of $\beta$-mixing coefficients under uniform ergodicity
Under uniform ergodicity, the mixing time is finite, and the $\beta$-mixing coefficients decays exponentially. In this setting, we show how to obtain a sample complexity upper bound for high-probability estimates on the individual $\beta$-mixing coefficients.
Theorem 3.2.
Let $s,n\in\mathbb{N}$ with $n>2s+1$, let $p\in\mathbb{R}_{+},p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any uniformly ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$,
  1. $(i)$
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 8\sqrt{\frac{1}{n-1}\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}},\\ \end{split}$
    where $t_{\mathsf{mix}}$ is the mixing time of $P$.
  2. $(ii)$
    Let $\varepsilon,\delta\in(0,1)$. For
    $n\geq 1+\frac{2^{8}}{\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{2}{\delta}\right\},$
@@ 2. $(ii)$
Using a truncated estimator with a cut-off determined from a uniform bound on the complexity terms $\|Q^{(s)}\|_{1-1/p}^{1-1/p}$, we can control the sequence of $\beta$-mixing coefficients with respect to the uniform norm modulo an additional logarithmic factor in the sample complexity.
Theorem 3.3.
Let $\varepsilon\in(0,1/4],\delta\in(0,1),p\in\mathbb{R}_{+},p\geq 1$ and $\mathcal{J}_{p}\in\mathbb{R}_{+}$. Let $n,S\in\mathbb{N}$, and for $s\in\mathbb{N}$, let $\widehat{\beta}_{[S]}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined as
$\widehat{\beta}_{[S]}(s)=\begin{cases}\widehat{\beta}(s)&\text{ for }s\leq S\\ 0&\text{ otherwise,}\end{cases}$
where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$ such that,
$\sup_{s\in\mathbb{N}}\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\leq\mathcal{J}_{p},$
when
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\varepsilon}\right\rceil\max\left\{\mathcal{J}_{p},18\log\left(\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\varepsilon}\right\rceil\right)\right\},$
with probability at least $1-\delta$, it holds that,
$\left\|\widehat{\beta}_{[S(n,\varepsilon,\delta,p,\mathcal{J}_{p})]}-\beta\right\|_{\infty}\leq\varepsilon,$
where the cut-off is set as
$S(n,\varepsilon,\delta,p,\mathcal{J}_{p})\doteq\frac{(n-1)\varepsilon^{2}}{2^{9}\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}}.$
@@
We note but do not pursue the fact that a more delicate bound on the sample complexity in Theorem 3.3 could be obtained, replacing a logarithm with a Lambert $W$ function.
In Section 4, we show how to obtain bounds on the entropy term in several natural settings.
@@ 3.4   Estimation of the average-mixing time
We move on to the estimation of the average-mixing time. Since the average distance to stationarity $d^{\sharp}=\beta$ does not generally enjoy sub-multiplicativity (it does not hold that $t^{\sharp}_{\mathsf{mix}}(\xi)$ can be bounded by $t^{\sharp}_{\mathsf{mix}}(1/4)$ and a polynomial of $\log\frac{1}{\xi}$), we formalize desired guarantees somewhat differently from Hsu et al. (2019); Wolfer and Kontorovich (2024). Indeed, for the relaxation time $t_{\mathsf{rel}}$, we were seeking an estimator $\widehat{t_{\mathsf{rel}}}$ to multiplicative error $\varepsilon$. Namely, the estimators in Hsu et al. (2019); Wolfer and Kontorovich (2024) guarantee that with high probability,
$\left|\frac{\widehat{t_{\mathsf{rel}}}}{t_{\mathsf{rel}}}-1\right|\leq\varepsilon.$
Here, we keep the distance to stationarity as a free parameter, and construct an estimator for the average-mixing time for proximity parameter $\xi$ up to multiplicative accuracy $\varepsilon$. Define for $\xi\in(0,1)$,
@@
where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (5).
Theorem 3.4.
Let $\varepsilon\in(0,1/4),\delta,\xi\in(0,1),p\in\mathbb{R}_{+}$ and $\mathcal{J}_{p,\xi}\in\mathbb{R}_{+}$. Let $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (6). For any uniformly ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$ such that,
$\sup\left\{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\}\leq\mathcal{J}_{p,\xi},$
when
$n\geq 1+\frac{2^{8}t_{\mathsf{mix}}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+\frac{p}{\log 2}\right)\mathcal{J}_{p,\frac{3\xi}{4}},9\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+1\right)\log\left(\frac{4t_{\mathsf{mix}}}{\delta}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil\right)\right\},$
@@
3.5   Extension to non-stationary Markov chains
Although results in this section are stated assuming a stationary start, it is possible to generalize them to non-stationary chains as follows. Let $\mu\in\mathcal{P}(\mathcal{X})$ absolutely continuous with respect to $\pi$. Inspired by the proof of Paulin (2015, Proposition 3.10), an application of Hölder’s inequality with $q,r\geq 1$ with $11/q+1/r=$ 1 yields,
3.6   Comparison with an alternative method
We point out the fact that for the special case $p=2$, by using the techniques in Paulin (2015), one can obtain an upper bound of the order (ignoring logarithmic factors) of
$\frac{t_{\mathsf{mix}}}{\varepsilon^{2}\xi^{2}}\mathcal{J}_{2,\xi}$
for the problem of estimating the average-mixing time in uniformly ergodic Markov chains. Indeed, the mixing time of the $s$-skipped pairs of observations is upper bounded by $t_{\mathsf{mix}}^{(s)}+s\leq\lceil t_{\mathsf{mix}}/s\rceil+s$. From Paulin (2015, Theorem 3.2, Proposition  3.4), it follows that
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]&\leq 2\lfloor(n-1)/s\rfloor\lceil t_{\mathsf{mix}}/s\rceil\pi(x),\\ \operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]&\leq 2\lfloor(n-1)/s\rfloor(\lceil t_{\mathsf{mix}}/s\rceil+s)\pi(x)P^{s}(x,x^{\prime}),\\ \end{split}$
which yields
$\begin{split}\mathbb{E}\left|\widehat{\beta}(s)-\beta(s)\right|\leq&\sqrt{2\frac{\lceil t_{\mathsf{mix}}/s\rceil+s}{\lfloor(n-1)/s\rfloor}}\left(\left\|Q^{(s)}\right\|_{1/2}^{1/2}+2\left\|\pi\right\|_{1/2}^{1/2}\right).\\ \end{split}$
Our method is superior in that it introduces a trade-off parameter $p$ to obtain an upper bound of order
$\frac{t_{\mathsf{mix}}}{\varepsilon^{2}\xi^{2}}\inf_{p\geq 1}\left\{p\mathcal{J}_{p,\xi}\right\}.$
This enables the use of more versatile bounds on the entropic term, which we explore in Section 4.

### v4

@@ 3.1  Estimation of $\beta$-mixing coefficients
We use the shortcut notation $N_{x}=N_{x}^{(1)}$ and $N_{xx^{\prime}}=N_{xx^{\prime}}^{(1)}$ to respectively denote number of visits to state $x$ and transitions from $x$ to $x^{\prime}$ on the original chain. For $s<n$, we define the following estimator $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ for $\beta(s)$,
@@
$\widehat{\beta}(s)\doteq\frac{1}{2\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|,$ (8)
and by convention, we set $\widehat{\beta}(s)\doteq 0$ for $s\geq n$.
3.1.1 Under general ergodicity condition
Under general—possibly non uniform—ergodicity, the mixing time may be infinite, and the $\beta$-mixing coefficients may decay at arbitrary slow rate. We first analyze the convergence properties of our estimator $\widehat{\beta}(s)$ under an arbitrary $\beta$-mixing rate, and will subsequently focus on exponential, sub-exponential and polynomial rates of convergence (refer to Definition 2.1).
Theorem 3.1 (Mean Absolute Deviation).
@@ 3.1.1 Under general ergodicity condition
Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+}$, with $p\geq 1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (8). For any $\pi$-stationary ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, it holds that
@@
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq 3\sqrt{\frac{\left(1/2+\mathcal{B}_{p}^{(s)}\right)\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
with
@@
$\mathcal{B}_{p}^{(s)}\doteq\sum_{t=0}^{\lfloor(n-1)/s\rfloor-1}\beta(st)^{1/p},\qquad\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|^{1-1/p}_{(1-1/p)/2},$
where for $s\in\mathbb{N}$, $\beta(s)$ follows Definition 4, for $x,x^{\prime}\in\mathcal{X}$, we wrote $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$, and for $q\in\mathbb{R}_{+}$,
$\left\|Q^{(s)}\right\|^{q}_{q}\doteq\sum_{(x,x^{\prime})\in\mathcal{X}^{2}}Q^{(s)}(x,x^{\prime})^{q}.$
Theorem 3.2 ($(\varepsilon,\delta)$-PAC Bound).
@@
Let $\varepsilon,\delta\in(0,1)$, $s,n\in\mathbb{N}$ with $n>2s+1$, $p\in\mathbb{R}_{+}$, with $p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (8). There exists a universal constant $C_{\mathsf{erg.}}\in\mathbb{R}_{+}$ such that for any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, when
$n\geq 1+C_{\mathsf{erg.}}\frac{\log(1/\delta)}{\varepsilon^{2}}\max\left\{\frac{s}{\varepsilon^{2}}\mathcal{B}_{p}^{(s)}\mathcal{J}^{(s)}_{p},t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right)\right\},\qquad\xi(\varepsilon,\delta)=\frac{\varepsilon^{2}\delta}{C_{\mathsf{erg.}}\log(1/\delta)},$
with probability at least $1-\delta$, it holds that
$\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon.$
In Theorem 3.1 and Theorem 3.2, we observe that the two quantities $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$, discussed below, control the estimation rate. In infinitely denumerable state spaces, for any $s\in\mathbb{N}$, the well-definition of the estimation rate therefore hinges on the existence of a $p\in[1,\infty)$ such that both $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$ are finite.
The $\beta$-mixing parameter $\mathcal{B}_{p}^{(s)}$. This parameter summarizes the $\beta$-mixing properties of the process and is controlled from above by the $\ell_{1/p}$ pseudo-norm of the sequence of the $\beta^{(s)}$-mixing coefficients,
$\mathcal{B}_{p}^{(s)}\leq\left\|\beta^{(s)}\right\|_{1/p}^{1/p}.$
It will be convenient to introduce the shorthand notation $\mathcal{B}_{p}\doteq\mathcal{B}_{p}^{(1)}$ to refer to the sum of the $\beta$-mixing coefficients. It is instructive to observe that since $\beta(s)\leq 1$ for any $s\in\mathbb{N}$, $\mathcal{B}_{p}^{(s)}$ increases together with $p$, and
$\lim_{p\to\infty}\mathcal{B}_{p}^{(s)}\asymp\frac{n}{s}.$
Additionally,
$r\leq s\implies\mathcal{B}_{p}^{(r)}\geq\mathcal{B}_{p}^{(s)}.$
In the following Lemma 3.1, we readily provide more explicit upper bounds on $\mathcal{B}_{p}^{(s)}$ under additional assumptions on the mixing rate.
Lemma 3.1.
Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$. For any $\pi$-stationary ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$, the following statements hold.
@@
[…]
@@
    $\begin{split}\mathcal{B}_{p}^{(s)}&\leq 1+\exp(-\beta_{1}s^{b}/p)+\frac{\Gamma\left(\frac{1}{b},s^{b}\beta_{1}/p\right)}{bs(\beta_{1}/p)^{1/b}},\end{split}$
    where $\Gamma$ is the (upper) incomplete Euler gamma function, given by $\Gamma(a,x)=\int_{x}^{\infty}t^{a-1}e^{-t}dt$.
@@
    $\begin{split}\mathcal{B}_{p}^{(s)}&\leq\frac{\beta_{0}^{1/p}+\zeta(b/p)\frac{\beta_{1}^{1/p}}{s^{b/p}}}{\lfloor(n-1)/s\rfloor},\end{split}$
@@
The entropic parameter $\mathcal{J}_{p}^{(s)}$. For $q<1$, since the $\ell_{q}$ quasi-norm and the $q$th order Rényi entropy $H_{q}$ of a distribution $\mu\in\mathcal{P}(\mathcal{X})$ are related by the expression
$\frac{1}{1-q}\log\left\|\mu\right\|_{q}^{q}=H_{q}(\mu),$
the quantity $\mathcal{J}_{p}^{(s)}$ pertains to an entropic property inherent in the Markov process. Observe that in contradistinction with $\mathcal{B}_{p}^{(s)}$, the quantity $\mathcal{J}_{p}^{(s)}$ decreases with $p$, leading to a trade-off in the choice of the parameter $p$. A necessary condition on $P$ for a rate to exist is readily given by
$\mathcal{J}_{\infty}^{(s)}\doteq\lim_{p\to\infty}\mathcal{J}_{p}^{(s)}=\left\|Q^{(s)}\right\|_{1/2}<\infty,$ (9)
that is, the tail of the distribution of stationary pairs of observations separated by a time $s$ must not decay more slowly than an inverse square polynomial. A more in-depth analysis of $\mathcal{J}_{p}^{(s)}$ is deferred to Section 4, where we show how to obtain upper bounds under natural assumptions.
3.1.2 Under uniform ergodicity
Under uniform ergodicity, the mixing time is finite and the $\beta$-mixing coefficients decay exponentially. In this setting, we show how to obtain a stronger sample complexity upper bound for high-probability estimates on the individual $\beta$-mixing coefficients. In particular, uniform ergodicity removes the need to consider the trade-off between mixing and entropic parameters mentioned in the previous section.
Theorem 3.3 (Mean Absolute Deviation and $(\varepsilon,\delta)$-PAC Bound).
Let $s,n\in\mathbb{N}$ with $n>2s+1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (8). There exists a universal constant $C_{\mathsf{u.erg.}}\leq 3\sqrt{2}$ such that for any $\pi$-stationary uniformly ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$,
[…]
where $t_{\mathsf{mix}}$ is the mixing time of $P$ and $\mathcal{J}_{\infty}^{(s)}$ is defined in (9).
Theorem 3.4 ($(\varepsilon,\delta)$-PAC Bound).
Let $\varepsilon,\delta\in(0,1),s,n\in\mathbb{N}$ with $n>2s+1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (8). There exists a universal constant $C_{\mathsf{u.erg.}}\leq 3\sqrt{2}$ such that for any $\pi$-stationary uniformly ergodic Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$, for
$n\geq 1+\frac{4(t_{\mathsf{mix}}+2s)}{\varepsilon^{2}}\max\left\{C_{\mathsf{u.erg.}}^{2}\mathcal{J}_{\infty}^{(s)},576\log\frac{2}{\delta}\right\},$
@@ 3.1.2 Under uniform ergodicity
Specializing Theorem 3.1 to obtain a PAC bound in the uniformly ergodic setting would instead yield an inferior PAC upper bound in
$n\geq\frac{C}{\varepsilon^{2}}\max\left\{\left(s+pt_{\mathsf{mix}}\right)\mathcal{J}_{p}^{(s)},(s+t_{\mathsf{mix}})\log\frac{1}{\delta}\right\},C>0.$
3.1.3 Comparison with prior results
In McDonald et al. (2015, Theorem 4), the authors establish the nearly parametric convergence rate for estimating $\beta$-mixing coefficients of continuously valued Markov processes
where $W_{0}$ is the Lambert function (Corless et al., 1996). They also posit that their findings may be considered conservative when applied to discrete-valued processes. Theorem 3.1 validates their remark by recovering the superior parametric rate $\mathcal{O}(\sqrt{1/n})$ under convergence of $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$. In the general ergodic setting, our method shares similarities with Khaleghi and Lugosi (2023), as both employ Rio’s covariance inequality Rio (1999, Corollary 1.1). However, compared to both Khaleghi and Lugosi (2023) and McDonald et al. (2015, Theorem 4), we distinguish ourselves by focusing on the discrete space setting and deriving explicit rates in terms of fine-grained properties of the transition operator. We also note that Grünewälder and Khaleghi (2026) address the estimation of $\beta$-mixing coefficients in the geometrically ergodic setting, while we consider a broader class of $\beta$-mixing processes, including sub-exponentially and algebraically $\beta$-mixing ones. Finally, under uniform ergodicity, instead of Rio’s covariance inequality, we rely on Paulin (2015) which yields a stronger bound on the MAD and sharper PAC bounds (refer to Remark 3.1).
@@ 3.2  Estimation of the average-mixing time
We now show how the sequence of estimators for the $\beta$-mixing coefficients, discussed in the previous section, translates into an estimator for the average-mixing time. Recall that the average distance to stationarity $\beta$ does not generally enjoy sub-multiplicativity. In turn it $t^{\sharp}_{\mathsf{mix}}(\xi)$ cannot generally be bounded by $t^{\sharp}_{\mathsf{mix}}(1/4)$ and a polynomial of $\log(1/\xi)$. Consequently, we formalize desired guarantees on our estimator differently from Hsu et al. (2019); Wolfer and Kontorovich (2024) for the relaxation time, where the error is defined multiplicatively with respect to the estimator $|\widehat{t_{\mathsf{rel}}}/t_{\mathsf{rel}}-1|\leq\varepsilon$. When estimating the average-mixing time, we treat the proximity $\xi$ to stationarity as a free parameter, and construct an estimator that targets a band $\xi(1\pm\varepsilon)$. Specifically, we define for $\xi\in(0,1)$,
@@
where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (8), and our next objective is to obtain with high probability an upper bound on the trajectory length such that
As illustrated on Figure 2, the interval above becomes thinner as $\varepsilon$ decays to zero. Our next lemma explains how to convert a sufficient trajectory length for high-probability control of the error of the sequence of estimators for the $\beta$-mixing coefficients into a sufficient trajectory length in order to estimate $t^{\sharp}_{\mathsf{mix}}(\Theta(\xi))$.
[…]
Figure 2: Average-mixing time $\xi(1\pm\varepsilon)$ band.
Lemma 3.2.
Let $\delta,\xi,\varepsilon\in(0,1)$ and $n\in\mathbb{N}$. Suppose that for any $s\in\mathbb{N}$, there exists an estimator $\widehat{\beta}(s)$ for $\beta(s)$ such that for any $\pi$-stationary ergodic Markov chain $X_{1},\dots,X_{n}$, when $n\geq n_{0}(P,s,\varepsilon,\delta)$ it holds with probability at least $1-\delta$ that
$\left|\widehat{\beta}(s)-\beta(s)\right|<\varepsilon.$
Let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (10). Then for
$n\geq\max_{1\leq s\leq t^{\sharp}_{\mathsf{mix}}(\xi(1-\varepsilon))}n_{0}\left(P,s,\xi\varepsilon,\frac{\delta}{2t^{\sharp}_{\mathsf{mix}}(\xi)}\right),$
@@
Theorem 3.5.
Let $\delta,\xi,\varepsilon\in(0,1)$, $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (10). There exists a universal constant $C\in\mathbb{R}_{+}$ such that the following holds. Let $X_{1},X_{2},\dots,X_{n}$ be a $\pi$-stationary uniformly ergodic Markov with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$. For any $p\in\mathbb{R}_{+}\cup\{\infty\}$, we write
$\mathcal{J}_{p,\xi}\doteq\sup\left\{\mathcal{J}_{p}^{(s)}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\}.$
  1. 1.
    When $P$ is ergodic, and
    $n\geq 1+C_{\mathsf{erg.}}^{2}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right)\frac{1}{\xi^{4}\varepsilon^{4}}\min_{p\geq 1}\left\{\mathcal{B}_{p}\mathcal{J}_{p,\xi(1-\varepsilon)}\right\}\log\frac{4t^{\sharp}_{\mathsf{mix}}(\xi)}{\delta},\qquad\xi(\varepsilon,\delta)=\frac{\varepsilon^{2}\delta}{C_{\mathsf{erg.}}\log(1/\delta)},$
  2. 2.
    or when $P$ is uniformly ergodic, and
    $n\geq 1+4t_{\mathsf{mix}}\frac{1+\left\lceil\log_{2}\frac{1}{\xi(1-\varepsilon)}\right\rceil}{\xi^{2}\varepsilon^{2}}\max\left\{C_{\mathsf{u.erg.}}^{2}\mathcal{J}_{\infty,\xi(1-\varepsilon)},576\log\frac{4t_{\mathsf{mix}}\left\lceil\log_{2}1/\xi\right\rceil}{\delta}\right\},$
then with probability at least $1-\delta$, it holds that
Proof.
The theorem follows from plugging Theorem 3.2 and Theorem 3.4 into Lemma 3.2. The theorem in the uniformly ergodic case follows from
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq t_{\mathsf{mix}}(\xi)\leq t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil.$
Corollary 3.1 (Confidence interval for $t^{\sharp}_{\mathsf{mix}}(\xi)$).
As a consequence of Theorem 3.5, we can construct a confidence interval for $t^{\sharp}_{\mathsf{mix}}(\xi)$. Denoting $n_{0}(\xi,\delta)$ the sample complexity in Theorem 3.5, when
$n\geq\max\left\{n_{0}\left(\frac{\xi}{1-\varepsilon},\delta/2\right),n_{0}\left(\frac{\xi}{1+\varepsilon},\delta/2\right)\right\},$
it holds with probability $1-\delta$ that
$t^{\sharp}_{\mathsf{mix}}(\xi)\in\left[\widehat{t_{\mathsf{mix}}}^{\sharp}\left(\frac{\xi}{1+\varepsilon}\right),\widehat{t_{\mathsf{mix}}}^{\sharp}\left(\frac{\xi}{1-\varepsilon}\right)\right].$
In the non uniformly ergodic setting, our method introduces a trade-off parameter $p$ to obtain an upper bound of order
$\widetilde{\mathcal{O}}\left(\frac{t^{\sharp}_{\mathsf{mix}}}{\xi^{4}}\inf_{p\geq 1}\left\{\mathcal{B}_{p}\mathcal{J}_{p,\xi}\right\}\right).$
This enables the use of versatile bounds on the entropic term, which we explore in Section 4. Currently, there are no related work to which we can compare the general results obtained in Section 3.2. We defer a comparison to the state of the art in the finite space setting to Section 4.3.